OpenStudy (anonymous):

simplifiy f(x)=x^2-1 if f(4+h)

OpenStudy (anonymous):

(4+h)(4+h)-1 h^2+8h+15

OpenStudy (anonymous):

insert (4+h) into f(x)=X^2-1 f(4+h)=f(x) so x=(4+h)

OpenStudy (anonymous):

got it miller

OpenStudy (anonymous):

got it : )